# Practical indistinguishability in a gene regulatory network inference problem, a case study

### v1

@@ 6.3 Family of regulatory models descriptions
$\text{[eud-1 mRNA]:}~\dfrac{dM_{e}}{dt}$ $=\alpha_{1}+k_{t_{e}}F_{1}(E,N,S)-k_{-m}M_{e},$ (23)
@@
$\text{[nhr-40 mRNA]:}~\dfrac{dM_{n}}{dt}$ $=\alpha_{2}+k_{t_{n}}F_{2}(E,N,S)-k_{-m}M_{n},$ (25)
@@
$\text{[sult-1 mRNA]:}~\dfrac{dM_{s}}{dt}$ $=\alpha_{3}+k_{t_{s}}F_{3}(E,N,S)-k_{-m}M_{s},$ (27)
@@
$\text{[nhr-40 mRNA]:}~\dfrac{dM_{n}}{dt}$ $=\alpha_{2}+k_{t_{n}}G_{2}(E,N)-k_{-m}M_{n},$ (29)
@@
$\text{[sult-1 mRNA]:}~\dfrac{dM_{s}}{dt}$ $=\alpha_{3}+k_{t_{s}}G_{3}(E,N)-k_{-m}M_{s},$ (31)
@@
$\text{[eud-1 mRNA]:}~\dfrac{dM_{e}}{dt}$ $=\alpha_{1}+k_{t_{e}}H_{1}(N,S)-k_{-m}M_{e},$ (34)
@@
$\text{[nhr-40 mRNA]:}~\dfrac{dM_{n}}{dt}$ $=\alpha_{2}+k_{t_{n}}H_{2}(N,S)-k_{-m}M_{n},$ (36)
@@ 6.4 Structural identifiability testing
As shown in Section 3, the model library contains a scaling symmetry (see Eqs. (1)-(6)). This scaling symmetry impacts Eqs. (23)-(28), Eqs. (29)-(32), Eqs. (34)-(37). We identified the scaling symmetry and found a reparameterization of the tested model using STRIKE-GOLDD [18]. To eliminate the scaling symmetry, we set $\xi=\dfrac{1}{k_{t}}$, which rescales the protein states and the half-max parameters by the translation rate, which must be included in the model and can never be zero. Therefore, we scaled the by the translation rate. By inspection, we concluded that the scaling symmetry would impact the entire library, given the conserved structure of the equations.
To ensure the reparameterized models were structurally identifiable, we attempted to screen the entire family using an independent algorithm [19]. We screened the identifiability of all kinetic parameters and initial conditions for the entire family of ODE models (including both the wild-type models and the knock-out models) using the same algorithm, and found that thousands of models were indeed structurally identifiable, but this approach was computationally infeasible for many of the highly nonlinear models. Many highly non-linear models did not finish in 24 hours on an HPC cluster, so we cannot conclude if they are structurally identifiable or not. It is known that some model structures can cause long algorithm run-times [19]. Of the thousands of models that finished within 24 hours, we found that the kinetic parameters were identifiable, and the initial conditions of the scaled protein states were identifiable if they appeared in at least one of the mRNA production terms in the model. It should be noted that this finding is based on testing a limited number of models. It is also important to note that the family of ODE models could contain a more complex scaling or transformation symmetry that is impacting structural identifiability that we failed to detect, even though we used multiple algorithms to assess structural identifiability [18, 19].
@@ 6.5 Parameter estimation
We performed parameter estimation across the model structures for each experimental data set using the L-BFGS-B algorithm initialized from 1,000 places in parameter space using Latin-Hypercube sampling(see Sections 6.3 and 6.4) [30]. As we did not have good prior estimates, we chose very wide lower and upper bounds for unscaled parameters, between 1x10-5 and 1x104+1, respectively. For parameters scaled by the translation rate (see Section 6.4), we allowed parameters to range from 1x10-9 and 1x108+1. The Latin-Hypercube samples were logarithmically rescaled between 1x10-4 and 1x104 and 1x10-8 and 1x108 for unscaled and scaled parameters, respectively. Smaller parameter ranges cause some parameters to hit the upper or lower bounds we set. Tuning optimization hyperparameters, such as the upper and lower bounds, for individual model structures was infeasible due to the large number of models. We assumed the data followed a multiplicative error model and minimized the associated negative log-likelihood function across all experiments. The negative log-likelihood function has the form
@@ 6.6 Model sets calculation
After fitting the models of interest (see Sections 6.2.1, 6.3, 6.4, 6.5), we built model sets, or combinations of features that are shared across the models that generate acceptable fits to the data. It was feasible to build the model sets associated with each of the knock-out experiments, given the relatively few models involved and the clear discrete jump in the cost function value (see Figure 5 A, middle and right). Building the model set associated with the wild-type experimental data was more involved.
@@
To find the model set associated with the wild-type experiment, we first binarized 21 model features of the 13,824 model structures we tested (see Figure 4D). We then used a decision tree to predict if the models gave an acceptable or unacceptable fit to the data based on the model features. We considered the models with the 1000 lowest cost function values to be acceptable fits and all other model fits to be unacceptable based on the behavior of the cost function value curve in Figure 4A. After 1000 models, the cost function value curve plateaus, then increases dramatically (see Figure 4A). We used K-Folds cross-validation and found an average accuracy of 94.8%. Note that this is higher than a null-hypothesis of all models do not produce an acceptable fit (92.8% accuracy).
@@
We explored up-sampling the minority class and found that it led to a marginal improvement in accuracy, but also led to overfitting and did not qualitatively change the results. To decide on the depth of the tree, we swept over the depth parameter of the decision tree, and found that accuracy was maximized when the maximum depth was between five and seven. We set the criterion to the Gini index or entropy. For simplicity, we set the maximum depth to five. Using this approach, we were able to detect a structural signal of the combination of model features that generate an acceptable fit to the data, but were not able to correctly classify all models. This is unsurprising, as the cost function value curve exhibits a rounded plateau (Figure 4A).

### v2

@@ 6 Methods & Technical Notes
Code associated with this project can be found at https://github.com/cefitzg/mouthform_code. Supercomputing code can be made available upon request.
@@ 6.3 Family of regulatory models descriptions
$\text{[eud-1 mRNA]:}~\dfrac{dM_{e}}{dt}$ $=\alpha_{1}+k_{t_{e}}F_{1}(E,N,S,K_{1},K_{4})-k_{-m}M_{e},$ (23)
@@
$\text{[nhr-40 mRNA]:}~\dfrac{dM_{n}}{dt}$ $=\alpha_{2}+k_{t_{n}}F_{2}(E,N,S,K_{2},K_{5})-k_{-m}M_{n},$ (25)
@@
$\text{[sult-1 mRNA]:}~\dfrac{dM_{s}}{dt}$ $=\alpha_{3}+k_{t_{s}}F_{3}(E,N,S,K_{3},K_{6})-k_{-m}M_{s},$ (27)
@@
$\text{[nhr-40 mRNA]:}~\dfrac{dM_{n}}{dt}$ $=\alpha_{2}+k_{t_{n}}G_{2}(E,N,K_{2},K_{5})-k_{-m}M_{n},$ (29)
@@
$\text{[sult-1 mRNA]:}~\dfrac{dM_{s}}{dt}$ $=\alpha_{3}+k_{t_{s}}G_{3}(E,N,K_{3},K_{6})-k_{-m}M_{s},$ (31)
@@
$\text{[eud-1 mRNA]:}~\dfrac{dM_{e}}{dt}$ $=\alpha_{1}+k_{t_{e}}H_{1}(N,S,K_{1},K_{4})-k_{-m}M_{e},$ (34)
@@
$\text{[nhr-40 mRNA]:}~\dfrac{dM_{n}}{dt}$ $=\alpha_{2}+k_{t_{n}}H_{2}(N,S,K_{3},K_{6})-k_{-m}M_{n},$ (36)
@@ 6.4 Structural identifiability testing
As shown in Section 3, the model library contains a scaling symmetry (see Eqs. (1)-(6)). This scaling symmetry impacts Eqs. (23)-(28), Eqs. (29)-(32), Eqs. (34)-(37). We identified the scaling symmetry and found a reparameterization of the tested model using STRIKE-GOLDD [27]. To eliminate the scaling symmetry, we set $\xi=\dfrac{1}{k_{t}}$, which rescales the protein states and the half-max parameters by the translation rate, which must be included in the model and can never be zero. Therefore, we scaled the proteins and half-max parameters by the translation rate. By inspection, we concluded that the scaling symmetry would impact the entire library, given the conserved structure of the equations.
To ensure the reparameterized models were structurally identifiable, we attempted to screen the entire family using an independent algorithm [28]. We screened the identifiability of all kinetic parameters and initial conditions for the entire family of ODE models (including both the wild-type models and the knock-out models) using the same algorithm, and found that thousands of models were indeed structurally identifiable, but this approach was computationally infeasible for many of the highly nonlinear models. Many highly non-linear models did not finish in 24 hours on an HPC cluster, so we cannot conclude if they are structurally identifiable or not. It is known that some model structures can cause long algorithm run-times [28]. Of the thousands of models that finished within 24 hours, we found that the kinetic parameters were identifiable, and the initial conditions of the scaled protein states were identifiable if they appeared in at least one of the mRNA production terms in the model. It is also important to note that the family of ODE models could contain a more complex scaling or transformation symmetry that is impacting structural identifiability that we failed to detect, even though we used multiple algorithms to assess structural identifiability [27, 28].
@@ 6.5 Parameter estimation
We performed parameter estimation across the model structures for each experimental data set using the L-BFGS-B algorithm initialized from 1,000 places in parameter space using Latin-Hypercube sampling(see Sections 6.3 and 6.4) [39]. As we did not have good prior estimates, we chose very wide lower and upper bounds for unscaled parameters, between 1x10-5 and 1x104+1, respectively. For parameters scaled by the translation rate (see Section 6.4), we allowed parameters to range from 1x10-9 and 1x108+1. The Latin-Hypercube samples were logarithmically rescaled between 1x10-4 and 1x104 and 1x10-8 and 1x108 for unscaled and scaled parameters, respectively. Smaller parameter ranges cause some parameters to hit the upper or lower bounds we set. Tuning optimization hyperparameters, such as the upper and lower bounds, for individual model structures was infeasible due to the large number of models. We assumed the data followed a multiplicative error model  [^2] Specifically, we assume $\bar{y}=(1+\epsilon)\bar{x}$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and $\bar{y}$ is the vector of observed data. and minimized the associated negative log-likelihood function across all experiments. The negative log-likelihood function has the form
@@ 6.6 Model sets calculation
After fitting the models of interest (see Sections 6.2.1, 6.3, 6.4, 6.5), we built model sets, or models with shared regulatory network features that generate acceptable fits to the data. It was feasible to build the model sets associated with each of the knock-out experiments by hand, given the relatively few models involved and the clear discrete jump in the cost function value (see Figure 5 A, middle and right). Building the model set associated with the wild-type experimental data was more involved.
@@
To find the model set associated with the wild-type experiment, we first binarized 21 model features of the 13,824 model structures we tested. We then used a decision tree to predict if the models gave an acceptable or unacceptable fit to the data based on the model features (see Figure 6). We considered the models with the 1000 lowest cost function values to be acceptable fits (Figure 4 A, pink) and all other model fits to be unacceptable (Figure 4 A, gray) based on the behavior of the cost function value curve in Figure 4A. After 1000 models, the cost function value curve plateaus, then increases dramatically (see Figure 4A). We used K-Fold cross-validation and found an average accuracy of 94.8%. Note that this is higher than a null-hypothesis of all models do not produce an acceptable fit (92.8% accuracy).
@@
We explored up-sampling the minority class and found that it led to a marginal improvement in accuracy, but also led to overfitting and did not qualitatively change the results. To decide on the depth of the tree, we swept over the depth parameter of the decision tree, and found that accuracy was maximized when the maximum depth was between five and seven. We ran the analysis using both Gini index and entropy as criteria and found the result did not change. For simplicity, we set the maximum depth to five. Using this approach, we were able to detect a structural signal of the combination of model features that generate an acceptable fit to the data, but were not able to correctly classify all models. This is unsurprising, as the cost function value curve exhibits a rounded plateau (Figure 4A).
7 Author Contributions
C.E.F., M.S.W., and N.M.M. designed the research; C.E.F. and N.M.M. oversaw the research; C.E.F., A.M., and V.A. performed research; C.E.F., S.R., A.M., and M.S.W. wrote the paper; All authors reviewed and edited the paper.
8 Competing Interests
The authors declare no competing interest.
9 Supplemental Figure
Figure 6: (Caption continued on next page.)
: We used a decision tree to classify models that produced an acceptable fit to the wild type experiment (Figure 4 A, pink) from those that produced an unacceptable fit (Figure 4 A, gray) based on 21 distinct structural features of the 13,814 models (note: optimization did not finish within the 4 hour wall time for 10 models and are excluded from the analysis). The binary structural features of the model we used during the classification included are E_in_Me, N_in_Me, S_in_Me, Pos_Me, Neg_Me, Mix_Me, Degree_Me, E_in_Mn, N_in_Mn, S_in_Mn, Pos_Mn, Neg_Mn, Mix_Mn, Degree_Mn, E_in_Ms, N_in_Ms, S_in_Ms, Pos_Ms, Neg_Ms, Mix_Ms, Degree_Ms which represent $\hat{E}$ regulates eud-1 mRNA production term, $\hat{N}$ regulates eud-1 mRNA production term, $\hat{S}$ regulates eud-1 mRNA production term, eud-1 mRNA is solely positively regulated, eud-1 mRNA is solely negative regulated, eud-1 mRNA is modeled by a production term involving only one protein with an exponent of one, $\hat{E}$ regulates nhr-40 mRNA production term, $\hat{N}$ regulates nhr-40 mRNA production term, $\hat{S}$ regulates nhr-40 mRNA production term, nhr-40 mRNA is solely positively regulated, nhr-40 mRNA is solely negative regulated, nhr-40 mRNA is both positively and negatively regulated, nhr-40 mRNA is modeled by a production term involving only one protein with an exponent of one, $\hat{E}$ regulates sult-1 mRNA production term, $\hat{N}$ regulates sult-1 mRNA production term, $\hat{S}$ regulates sult-1 mRNA production term, sult-1 mRNA is solely positively regulated, sult-1 mRNA is solely negative regulated, sult-1 mRNA is both positively and negatively regulated, sult-1 mRNA is modeled by a production term involving only one protein with an exponent of one, respectively. The value 0 corresponds to FALSE and the value 1 corresponds to TRUE for each of the binary structural features of the model. ‘Gini’, ‘samples’, and ‘value’ are the Gini impurity value (a measure of split quality), the number of model structures that reach each node of the tree, and how samples are split by class (the first index corresponds to the unacceptable class and the second index corresponds to the acceptable class), respectively.